# Privacy-Preserving Neural Graph Databases

### 25 Dec 2023 v1

@@ 1. Introduction
Graph databases (graph DBs) play a crucial role in storing, organizing, and retrieving structured relational information and are widely applied in downstream applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. Unfortunately, traditional graph databases often suffer from the limitation of graph incompleteness, which is a prevalent issue in real-world knowledge graphs, like Freebase (Mintz et al., 2009; Bollacker et al., 2008). This incompleteness leads to the exclusion of relevant results, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013). To solve the drawbacks, extending the concept of Graph DBs, neural graph databases (NGDBs) combine the flexibility of graph data models with the computational capabilities of neural networks, enabling efficient representation, storage, and analysis of interconnected data (Ren et al., 2023; Hamilton et al., 2018).
@@
Figure 1. Structure of Neural Graph Databases (NGDBs). NGDBs can retrieve answers given complex logical queries on incomplete graphs. However, the privacy problem arises due to the storage of graph information in embedding space and the presence of a query engine that facilitates complex logical query answering. To illustrate the issue, consider an example where an attacker attempts to infer private information about Hinton’s living place in the NGDBs. Direct querying privacy can be easily detected by privacy risk detection, however, attackers can leverage the combination of multiple queries to retrieve desired privacy.
@@
NGDBs provide unified storage for diverse entities in an embedding space. By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020). As shown in Figure 1, logical expressions are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. The logical forms of $q$ include terms like $Win(V,Turing\ Award)$ and logical operators like conjunction $\wedge$, representing that find where the Turing Award winner who was born in 1960 lived. Query encoding methods are commonly used in Complex Query Answering (CQA). These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query embedding and entity embedding. For example, GQE (Hamilton et al., 2018) and query2box (Ren et al., 2020) encode queries as vector and hyper-rectangle, respectively.
@@
While NGDBs have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge graphs and enriches the retrieval information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. As shown in Figure 1, the living place information of an individual, in this case, Hinton as an example, is considered private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place. For example, by querying NGDBs with where Turing Award winners born before 1950 and after 1940 lived, where LeCun’s collaborator lived, etc. The intersection of these queries still presents a significant likelihood of exposing the living place of Turing Award winner Hinton.

### 19 Jan 2024 v2

@@ 1. Introduction
Graph databases (graph DBs) play a crucial role in storing, organizing, and retrieving structured relational information and are widely applied in downstream applications, such as recommender systems (Dong, 2018), fraud detection (Pourhabibi et al., 2020), etc. Unfortunately, traditional graph databases often suffer from the limitation of graph incompleteness, which is a prevalent issue in real-world knowledge graphs, like Freebase (Mintz et al., 2009; Bollacker et al., 2008). This incompleteness leads to the exclusion of relevant results, as the graph database may not capture all the necessary relationships and connections between entities by traversing (Bordes et al., 2013; Galárraga et al., 2013).
To solve the drawbacks, extending the concept of Graph DBs, Ren et al. propose neural graph databases (NGDBs) which combine the flexibility of graph data models with the computational capabilities of neural networks, enabling efficient representation, storage, and analysis of interconnected data (Ren et al., 2023).
@@
Figure 1. Privacy risks of Neural Graph Databases (NGDBs) facing malicious queries. NGDBs can retrieve answers given complex logical queries on incomplete graphs (Ren et al., 2023). However, the privacy problem arises due to the storage of graph information in embedding space and the presence of a query engine that facilitates complex logical query answering. To illustrate the issue, consider an example where an attacker attempts to infer private information about Hinton’s living place in the NGDBs. Direct querying privacy can be easily detected by privacy risk detection, however, attackers can leverage the combination of multiple queries to retrieve desired privacy.
@@
As shown in figure 1, NGDBs are designed with two modules: neural graph storage providing unified storage for diverse entities in an embedding space and neural query engine searching answers to input complex queries from the unified storage. By utilizing graph-based neural network techniques, these databases empower expressive and intelligent querying, inference, and knowledge discovery. This capability, known as complex query answering (CQA), aims to identify answers that satisfy given logical expressions (Hamilton et al., 2018; Ren et al., 2020). Logical expressions are often defined in predicate logic forms with relation projection operations, existential quantifiers $\exists$, logical conjunctions $\wedge$, disjunctions $\vee$, etc. The logical forms of $q$ include terms like $Win(V,Turing\ Award)$ and logical operators like conjunction $\wedge$, representing that find where the Turing Award winner who was born in 1960 lived. Query encoding methods are commonly used in Complex Query Answering (CQA). These methods involve parameterizing entities, relations, and logical operators, encoding both queries and entities into the same embedding space simultaneously, and retrieving answers according to the similarity between the query embedding and entity embedding. For example, GQE (Hamilton et al., 2018) and query2box (Ren et al., 2020) encode queries as vector and hyper-rectangle, respectively.
@@
While NGDBs have demonstrated significant achievements in various domains, they do face unique privacy challenges in comparison to traditional graph DBs. One notable risk arises from their generalization ability. Although generalization efficiently handles incomplete knowledge graphs and enriches the retrieval information, it also enables attackers to infer sensitive information from NGDBs by leveraging the composition of multiple complex queries. Figure 1 demonstrates an example of privacy attacks via malicious complex queries: the living place information of an individual, in this case, Hinton as an example, is considered private. Even though Hinton’s residence is omitted during the construction of the knowledge graph, or if direct privacy queries are restricted, a malicious attacker can still infer this sensitive information without directly querying Hinton’s living place. For example, by querying NGDBs with where Turing Award winners born before 1950 and after 1940 lived, where LeCun’s collaborator lived, etc. The intersection of these queries still presents a significant likelihood of exposing the living place of Turing Award winner Hinton.